OpenStudy (anonymous):

haloke does 176 joules of work lifting himself 0.300 meters. what is haloke's mass?

OpenStudy (anonymous):

The minimum force that Haloke must exert on himself to lift himself is equal to his weight. If he has mass m, then his weight will be: \[F_{g}=mg\]If he lifts himself a distance d, then the work he does is:\[W=Fd \sin \theta=F_{g}d=mgd\]It is given that W=176 J and d = 0.3 meters; g = 9.81 m/s^2. Now you can solve for m. ALTERNATE SOLUTION: The work done is equal in magnitude to the change in potential energy (assuming no energy is lost). Thus: \[W=\Delta PE = mg \Delta h\]Again, you are given the values for W and h, so you can easily find the mass.

OpenStudy (anonymous):

BY THE CONSERVATION OF ENERGY: WORKDONE=CHANGE IN POTENTIAL ENERGY 176 =Uf-Ui 176=mgh m=176/(10*.3)=58.66kg